OpenStudy (anonymous):

Given that f(x) = x² + 4x, evaluate f(-2).

OpenStudy (lifeisadangerousgame):

f(-2) means that you plug in -2 for each 'x' that you see. So: -2^2 + 4(-2) = ?

OpenStudy (anonymous):

f(-2) = (-2)^2 + 4(-2) f(-2) = 4 - 8 f(-2) = -4 You just plug the -2 in for the x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so the answer would be −12